40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 241 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  5208   Fri Aug 12 15:34:16 2011 NicoleSummarySUSBode Plot for TT Suspension

Quote:

Here is my bode plot comparing the flexibly-supported and rigidly-supported EDCs (both with no bar)

It seems as if the rigidly-supported EDC has better isolation below 10 Hz (the mathematically-determined Matlab model predicted this...that for the same magnet strength, the rigid system would have a lower Q than the flexible system). Above 10 Hz (the resonance for the flexibly-supported EDCs seem to be at 9.8 Hz) , we can see that the flexibly-supported EDC has slightly better isolation? I may need to take additional measurements of the transfer function of the flexibly-supported EDC (20 Hz to 100 Hz?)  to hopefully get a less-noisy transfer function at higher frequencies. The isolation does not appear to be that much better in the noisy region (above 20Hz). This may be because of the noise (possibly from the electromagnetic field from the shaker interfering with the magnets in the TT?). There is a 3rd resonance peak at about 22 Hz. I'm not sure what causes this peak...I want to confirm it with an FFT measurement of the flexibly-supported EDC (20 Hz to 40 Hz?)

 

 

EXPrigidvsflex.jpg

 

 

 Since the last post, I have found from the Characterization of TT data (from Jenne) that the resonant frequency of the cantilever springs for TT #4 (the model I am using) have a resonant frequency at 22 Hz. They are in fact inducing the 3rd resonance peak.

 

Here is a bode plot (CORRECTLY SCALED) comparing the rigidly-supported EDCs (model and experimental transfer functions)

RIGIDexp_model.jpg

 

Here is a bode plot comparing the flexibly-supported EDCs (model and experimental transfer functions). I have been working on this graph for FOREVER and with the set parameters, this is is close as I can get it (I've been mixing and matching parameters for well over an hour > <). I think that experimentally, the TTs have better isolation than the model because they have additional damping properties (i.e. cantilever blades that cause resonance peak at 22 Hz). Also, there may be a slight deviation because my model assumes that all four EDCs are a single EDC.

flexmodcomp.jpg

  5207   Fri Aug 12 15:16:56 2011 jamieUpdateSUStoday's SUS overview

Here's an update of the suspensions, after yesterdays in-vacuum work and OSEM tweaking:

  • PRM and ETMY are completely messed up.  The spectra are so bad that I'm not going to bother posting anything.   ETMY has extreme sensor voltages that indicate that it's maybe stuck to one of the OSEMS.  PRM voltages look nominal, so I have no idea what's going on there.
  • ITMY is much improved, but it could still use some work
  • SRM is a little worse than what it was yesterday, but we've done a lot of work on the ITMY table, such as moving ITMY suspension and rebalancing the table.
  • BS looks for some reason slightly better than it did yesterday
TM   M cond(B)
SRM SRM.png

      pit     yaw     pos     side    butt
UL    0.828   1.041   1.142  -0.135   1.057 
UR    1.061  -0.959   1.081  -0.063  -1.058 
LR   -0.939  -0.956   0.858  -0.036   0.849 
LL   -1.172   1.044   0.919  -0.108  -1.035 
SD    0.196  -0.024   1.861   1.000   0.043

4.20951
ITMY  ITMY.png       pit     yaw     pos     side    butt
UL    1.141   0.177   1.193  -0.058   0.922 
UR    0.052  -1.823   0.766  -0.031  -0.974 
LR   -1.948  -0.082   0.807  -0.013   1.147 
LL   -0.859   1.918   1.234  -0.040  -0.957 
SD   -1.916   2.178   3.558   1.000   0.635 
7.70705
BS  BS.png       pit     yaw     pos     side    butt
UL    1.589   0.694   0.182   0.302   1.042 
UR    0.157  -1.306   1.842  -0.176  -0.963 
LR   -1.843  -0.322   1.818  -0.213   0.957 
LL   -0.411   1.678   0.158   0.265  -1.038 
SD    0.754   0.298  -3.142   1.000   0.053
6.12779

 

  5206   Fri Aug 12 14:15:07 2011 NicoleSummarySUSBode Plot for TT Suspension

Here is my bode plot comparing the flexibly-supported and rigidly-supported EDCs (both with no bar)

It seems as if the rigidly-supported EDC has better isolation below 10 Hz (the mathematically-determined Matlab model predicted this...that for the same magnet strength, the rigid system would have a lower Q than the flexible system). Above 10 Hz (the resonance for the flexibly-supported EDCs seem to be at 9.8 Hz) , we can see that the flexibly-supported EDC has slightly better isolation? I may need to take additional measurements of the transfer function of the flexibly-supported EDC (20 Hz to 100 Hz?)  to hopefully get a less-noisy transfer function at higher frequencies. The isolation does not appear to be that much better in the noisy region (above 20Hz). This may be because of the noise (possibly from the electromagnetic field from the shaker interfering with the magnets in the TT?). There is a 3rd resonance peak at about 22 Hz. I'm not sure what causes this peak...I want to confirm it with an FFT measurement of the flexibly-supported EDC (20 Hz to 40 Hz?)

 

 

EXPrigidvsflex.jpg

 

 

  5205   Fri Aug 12 11:07:50 2011 NicoleSummarySUSMore TT Shaking Completed This Morning

This morning (about 10am to 11am), I have collected additional transfer function measurements for the T.T. suspension. I have finished taking my measurements. The SR785 has been returned to its place next the the seismometer racks.

 The data has been backed up onto the cit40m computer

  5204   Fri Aug 12 04:11:13 2011 kiwamuUpdateSUSfree swinging

Excited all optics - -
Fri Aug 12 03:34:12 PDT 2011
997180467

  5203   Fri Aug 12 03:52:51 2011 kiwamuUpdateSUSPRM OSEM adjustment

[Suresh / Kiwamu]

 We tried adjusting the OSEMs on PRM, but we didn't complete it due to a malfunction on the coils.

The UL and LL coils are not working correctly, the forces are weak.

Tomorrow we will look into the satellite box, which is one of the suspects.

 

 During the adjustment we found that the POS excitation force was unequal in each sensor.

At the beginning we thought it's because of the difference of the sensitivity in each OSEM due to the bad OSEM orientations.

However it turned out that it comes from the actual force imbalance on each coil.

We checked the force of each coil by putting an offset (-2000 cnts) in each output digital filter and looked at the OSEM signals in time series.

The UL and LL coils are too weak and the responses are almost buried in the noise of the OSEMs in time series.

We briefly checked some analog electronics and found the DAC, AI board and deWhitening board were healthy.

We were able to see the right amount of voltage from the monitor pin on the front panel of the coil driver.

So something downstream are suspicious, including the satellite box, feedthrough and coils.

- - -

Although the coil issue, it could be worth trying to check the input matrix.

  5202   Fri Aug 12 03:49:45 2011 JennySummaryPSLNPRO PDH-Locked to Ref Cav

DMass and I locked the NPRO laser (Model M126-1064-700, S/N 238) on the AP table to the reference cavity on the PSL table using the PDH locking setup shown in the block diagram below (the part with the blue background):

 

LIGO_block_diagram_2.png

 

A Marconi IFR 2023A signal generator outputs a sine wave at 230 kHz and 13 dBm, which is split. One output of the splitter drives the laser PZT while the other is sent to a 7dBm mixer. Also sent to the mixer is the output of a photodiode that is detecting the reflected power from off the cavity. (A DC block is used so that only RF signal from the PD is sent to the mixer). The output of the mixer goes through an SR560 low-noise preamp, which is set to act as a low pass filter with a gain of 5 and a pole at 30 kHz. That error signal is then sent to the –B port of the LB1005 PDH servo, which has the following settings: PI corner at 10kHz, LF gain limit of 50 dB, and gain of 2.7 (1.74 corresponds to a decade, so the signal is multiplied by 35). The output signal from the LB1005 is added to the 230 kHz dither using another SR560 preamp, and the sum of the signals drive the PZT.

 

I am monitoring the transmission through the cavity on a digital oscilloscope (not shown in the diagram) and with a camera connected to a TV monitor. I sweep the NPRO laser temperature set point manually until the 0,0 mode of the carrier frequency resonates in the cavity and is visible on the monitor. Then I close the loop and turn on the integrator on the LB1005.

 

The laser locks to the cavity both when the error signal is sent into the A port and when it is sent into the –B port of the PDH servo. I determined that –B is the right sign by comparing the transmission through the cavity on the oscilloscope for both ways.

 

When using the A port, the transmission when it was locked swept from ~50 to ~200 mV (over ~10 second intervals) but had large high frequency fluctuations of around +/- 50 mV. Looking at the error signal on the oscilloscope as well, the RMS fluctuations of the error signal were at best ~40 mV peak to peak, which was at a gain of 2.9 on the LB1005.

 

Using the –B port yielded a transmission that swept from 50 to 250 mV but had smaller high frequency fluctuations of around +/- 20 mV. The error signal RMS was at best 10mV peak to peak, which was at a gain of 2.7. (Although over the course of 10 minutes the gain for which the error signal RMS was smallest would drift up or down by ~0.1).

 

 

The open loop error signal peak-to-peak voltage was 180 mV, which is more than an order of magnitude larger than the RMS error signal fluctuations when the loop is closed, indicating that it is staying in the range in which the response is linear.

openlooperror.jpg

 

In the above plot the transmission signal is offset by 0.1 V for clarity.

Below is the closed loop error signal. The inset plot shows the signal viewed over a 1.6 ms time period. You can see ~60 microsecond fluctuations in the signal (~17 kHz)

closedlooperror.jpg

The system remained locked for ~45 minutes, and may have stayed locked for much longer, but I stopped it by opening the loop and turning off the function generator. Below is a picture of the transmitted light showing up on a monitor, the electronics I'm using, and a semi-ridiculous mess of wires.

 

IMG_3034.JPG

 

I determined that it’s not dangerous to leave the system locked and leave for a while. The maximum voltage that the SR560 will output to the PZT is 10Vpp. This means that it will not drive the PZT at more than +/-5 V DC. At low modulation rates, the PZT can take a voltage on the order of 30 Vpp, according to the Lightwave Series 125-126 user’s manual, so the control signal will not push the PZT too hard such that it’s harmful to the laser.

 

 

  5201   Fri Aug 12 00:18:30 2011 Ishwita, ManuelUpdatePEMCoherence of Guralp1 and STS2(Bacardi, Serial NR 100151)

 

 We moved the seismometer STS2(Bacardi, Serial NR 100151) as we told in this Elog Entry, so the distance between Guralp1 and STS2 is 31.1m. Following is the coherence plot for this case:

coher_gur1_sts1_31m.png

then we also moved the Guralp1 under the BS and plugged it with the Guralp2 cable (at 7:35pm PDT), so now the distance between the two seismometers is 38.5m. Following is the coherence plot for this case:

coher_gur1_sts1_38.png

  5200   Thu Aug 11 19:14:22 2011 Ishwita , ManuelUpdatePEMSeismometer STS2(Bacardi, Serial NR 100151) moved near ETMX

We moved the STS2(Bacardi, Serial NR 100151) to his new location and laid his cable from rack 1X7 to ETMX. The seismometer was below the mode cleaner vacuum tube before.

Now, (since 6:05pm PDT) its placed near the ETMX.

cable_to_etmx.pngsts2_etmx.png

 

  5199   Thu Aug 11 19:01:45 2011 SureshUpdateSUSITMX given a kick to start it freeswing' but to no avail.

We started an ITMX freeswing run at this time

Thu Aug 11 18:58:59 PDT 2011
997149554

 But the optic did not repond to the kick.  It is possible that the earthquake stops are close to the face and/or rear of the optic and prevent it from oscillating.  We will check again and see what is up in a few hours.

 

 

  5198   Thu Aug 11 18:30:40 2011 KojiUpdateSUSITMY OSEM adjustment

[Jamie, Koji]

ITMX OSEMs were adjusted so as to have the right DC numbers and the more uniform response to POS excitation.

It is waiting for the free-swinging test.

- ITMX was moved from its position to the north side of the table.

- The table was rebalanced.

- We found that the output of the LR OSEM has an excess noise compared with the other OSEMs.
We tried to swap the LR and SD OSEMs, but the SD OSEM (placed at the LR magnet) showed
the same excess noise at around 10-50Hz.

- We found that one of the EQ stops was touching the mirror. By removing this friction, all of the OSEMs
come to show similar power spectra. Good!

 - Then we started to use LOCKIN technique to measure the sensitivity of the OSEMs to the POS excitation.

Originally the response of the OSEMs was as follows
UL 3.4 UR 4.3 
LL 0    LR 2.5   

After the adjustment of the DC values, final values became as follows
UL 3.9 UR 4.4
LL 3.9  LR 3.2

- We decided to close the light door.

  5197   Thu Aug 11 16:21:16 2011 Ishwita , ManuelUpdatePEMPower spectra and Coherence of Guralps and STS2

 

 Following is the power spectrum plot (with corrected calibration [see here]) of seismometers Guralp1 and STS2(Bacardi, Serial NR 100151):

Pw_gur1_sts1_correct.png

 

 The seismometers are placed approximately below the center of the mode cleaner vacuum tube.

  5196   Thu Aug 11 16:15:59 2011 Ishwita, ManuelUpdatePEMCalibration of Guralp and STS2

 

Finally, we have found the correct calibration of Guralp and STS2 seismometers.

ADC: 216counts = 20V Hence, calibration of ADC is 3.2768e+03 counts/V.


 

GURALP

Sensitivity of seismometer = 800 V/ms-1

Gain of the guralp breakout box (reference elog entry) = 20

Calibration = 3.2768e+03 counts/V x 800  V/ms-1 x 20 = 52428800 counts/ms-1 -----> 1.9073e-08 ms-1/count


STS

Sensitivity = 1500 V/ms-1

Gain of the STS electronic breakout box = 10

Calibration = 3.2768e+03 counts/V x 1500  V/ms-1 x 10 = 49152000 counts/ms-1 -----> 2.0345e-08 ms-1/count

 

  5195   Thu Aug 11 16:09:05 2011 NicoleSummarySUSBode Plot for TT Suspension

All of my plots have already taken into account the calibration of the photosensor (V/mm ratio)

Here is a bode plot generated for the transfer function measurements we obtained last night/this morning. This is a bode plot for the fully-assembled T.T. (with flexibly-supported dampers and bottom bar). I will continue to upload bode plots (editing this post) as I finish them but for now I will go to sleep and come back later on today.

 

flexwithbase.jpg

Here is a bode plot comparing the no eddy-current damper case with and without the bar that we suspected to induce some non-uniform damping. We have limited data on the NO EDC, no bar measurements (sine swept data from 7 Hz to 50 Hz) and FFT data from 0 Hz to 12.5 Hz because we did not want to induce too much movement in the mirror (didn't want to break the mirror).  This plot shows that there is not much difference in the transfer functions of the TT (no EDC) with and without the bar.

stage1.jpg.jpg

From FFT measurements of  the no eddy-current damper case without the bar (800 data points, integrated 10 times) we can define the resonance peak of the TT mirror (although there are still damping effects from the cantilever blades).

The largest resonance peak occurs at about 1.94 Hz. The response (magnitude) is 230.

The second-largest resonance peak occurs at about 1.67 Hz. The response (magnitude) is 153. This second resonance peak may be due to pitch motion coupling (this is caused by the fact that the clamping attaching the mirror to the wires occurs above the mirror's center of mass, leading to inevitable linear and pitch coupling).

 

Here is a bode plot of the EDC without the bar. It seems very similar to the bode plot with the bar

FULL_BODE.jpg

 

 Here is a bode plot of the rigidly-supported EDC, without bar. I need to do a comparison plot of the rigid and flexibly-supported EDCs (without bar)

 

 FULL_BODErigid.jpg

 

Attachment 1: flexwithbase.jpg
flexwithbase.jpg
Attachment 3: stage1.jpg
stage1.jpg
  5194   Thu Aug 11 16:07:37 2011 steveUpdateSUSETMY rack cables strain releived
Attachment 1: P1080159.JPG
P1080159.JPG
Attachment 2: P1080160.JPG
P1080160.JPG
  5193   Thu Aug 11 15:31:43 2011 steveUpdateGeneralmalfunctioning crane at ETMY

Quote:

The horizontal trolley drive stopped working  at the east end this morning. It is working intermittently. In the worst case we can take the door off with the manual -Genie- lift.

I'm working with Konecrane to solve  the wormgear drive problem.

 Gear box found, it's  lead time one week at $825    The crane may be functional round August 25

  5192   Thu Aug 11 14:32:12 2011 KojiSummarySUSPhotosensor Head Calibration Curve for TT Frame

The entry was quite confusing owing to many misleading wordings.

- The PS2 should be calibrated "as is". (i.e. should be calibrated with the frame)

- The previous calibrations with the highly reflective surface were 0.32V/mm and 0.26V/mm, respectively.
  This time you have 0.10V/mm (with an undescribed surface). The ratio is not 32 but 3.2.

- The DC output of PS2 on the shaking setup was 2.5V. The DC output seen in the plot is 3.5V-ish.
This suggests the possibiliteies:
1) The surface has slightly higher reflectivity than the frame
2) The estimation of the distance between the frame and the PS2 during the TF measurement was not accurate.

- The word "DC coupling level" is misleading. I guess you mean the DC value of the vbration isolation transfer function
  of the suspension.

Quote:

I have re-calibrated the photosensor I used to measure the displacements of the TT frame (what I call "Photosensor 2").

As before, the linear region is about 15.2mm to 25.4mm. It is characterized by the slope -0.0996 V/mm (-0.1 V/mm). Recall that photosensor 1 (used to measure mirror displacements) has a calibration slope of -3.2V/mm. The ratio of the two slopes (3.2/0.1 = 32). We should thus expect the DC coupling level to be 32? This is not what we have for the DC coupling levels in our data (2.5 for flexibly-supported, fully-assembled TT (with EDC, with bar), 4.2 for EDC without bar, 3.2 for rigid EDC without bar, 3.2 for no EDC, with bar, 3.2 for no EDC without bar) . I think I may need to do my calibration plot for the photosensor at the frame?

ps2frame.jpg

 

  5191   Thu Aug 11 14:22:00 2011 NicoleSummarySUSPhotosensor Head Calibration Curve for TT Frame

Quote:

I have re-calibrated the photosensor I used to measure the displacements of the TT frame (what I call "Photosensor 2").

As before, the linear region is about 15.2mm to 25.4mm. It is characterized by the slope -0.0996 V/mm (-0.1 V/mm). Recall that photosensor 1 (used to measure mirror displacements) has a calibration slope of -3.2V/mm. The ratio of the two slopes (3.2/0.1 = 32). We should thus expect the DC coupling level to be 32? This is not what we have for the DC coupling levels in our data (2.5 for flexibly-supported, fully-assembled TT (with EDC, with bar), 4.2 for EDC without bar, 3.2 for rigid EDC without bar, 3.2 for no EDC, with bar, 3.2 for no EDC without bar) . I think I may need to do my calibration plot for the photosensor at the frame?

ps2frame.jpg

I have redone the voltage versus displacement measurements for calibrating "Photosensor 2" (the photosensor measuring the motions of the TT frame). This time, I calibrated the photosensor in the exact position it was in during the experimental excitation ( with respect to the frame ). I have determined the linear region to be 15.2mm to 22.9mm (in my earlier post today, when I calibrated the photosensor for another location on the frame, I determined the linear region to be 15.2mm to 25.4mm). This time, the slope was -0.92 V/mm (instead of -0.1 V/mm).

 

This means that the calibration ratio for photosensor 1 (measuring mirror displacements) and photoensor 2 (measuring frame displacements) is 34.86.

 

Since this "unity" value should be 34.86 for my transfer function magnitude plots (instead of the ~3 value I have), do I need to scale my data? It is strange that it differs by an order of magnitude...

  5190   Thu Aug 11 13:41:36 2011 Ishwita, ManuelUpdatePEMCoherence of Guralp1 and STS2(Bacardi, Serial NR 100151)

Following is the coherence plot obtained when Guralp1 and STS2(Bacardi, Serial NR 100151) are placed very close to each other (but they aren't touching each other):

coher_gur1_sts1_near.png

The seismometers were placed as shown in the picture below:

coherence_pic.JPG

They are placed below the center of the mode cleaner vacuum tube.

  5189   Thu Aug 11 12:54:06 2011 kiwamuUpdateIOOMC spot positions

The spot positions on the MC mirrors were adjusted by steering the MC mirrors, resulting in 1 mm off-centering on each optic.

DONE.

 

(Requirement cleared)

One of the requirements in aligning the MC mirrors is the differential spot positions in MC1 and MC3.

It determines the beam angle after the beam exists from MC, and if it's bigger than 3 mm then the beam will be possibly clipped by the Faraday (#4674).

The measured differential spot positions on MC1 and MC3 are : PIT = 0.17 mm and YAW = 1.9 mm, so they are fine.

 

(Measurement and Results)

 Suresh and I aligned the MC cavity's eigen axis by using MCASS and steering the MC mirrors.

Most of the alignment was done manually by changing the DC biases

because we failed to invert the output matrix and hence unable to activate the MCASS servo (#5167).

Then I ran Valera's script to measure the amount of the off-centering (#4355), but it gave me many error messages associated with EPICS.

So a new script newsensedecenter.csh, which is based on tdsavg instead of ezcaread, was made to avoid these error messages.

 

The resultant plot is attached. The y-axis is calibrated into the amount of the off-centering in mm.

In the plot each curve experiences one bump, which is due to the intentional coil imbalance to calibrate the data from cnts to mm (#4355).

The dashed lines are the estimated amount of off-centering.

For the definition of the signs, I followed Koji's coordinate (#2864) where the UL OSEM is always in minus side.

    Feb 26 2011      May 08 2011 Aug 2 2011 [NEW!!] Aug 10 2011 (in air)
MC1 pit [mm]   1.6   1.9  1.93 -0.858
MC2 pit [mm]   6.4   9.0 9.03 -0.844
MC3 pit [mm]   1.4   2.0 2.01 -1.03
MC1 yaw [mm]   -1.5   -1.7 -1.72 -0.847
MC2 yaw [mm]   1.0   0.2 0.178 0.582
MC3 yaw [mm]   -1.3   -1.9 -1.87 -1.06

 

Quote from #5182

After the beam spots on MC1 and MC3 were close to the actuation nodes (<1mm away)

Attachment 1: MCoffcenter.png
MCoffcenter.png
  5188   Thu Aug 11 12:31:39 2011 NicoleSummarySUSPhotosensor Head Calibration Curve for TT Frame

I have re-calibrated the photosensor I used to measure the displacements of the TT frame (what I call "Photosensor 2").

As before, the linear region is about 15.2mm to 25.4mm. It is characterized by the slope -0.0996 V/mm (-0.1 V/mm). Recall that photosensor 1 (used to measure mirror displacements) has a calibration slope of -3.2V/mm. The ratio of the two slopes (3.2/0.1 = 32). We should thus expect the DC coupling level to be 32? This is not what we have for the DC coupling levels in our data (2.5 for flexibly-supported, fully-assembled TT (with EDC, with bar), 4.2 for EDC without bar, 3.2 for rigid EDC without bar, 3.2 for no EDC, with bar, 3.2 for no EDC without bar) . I think I may need to do my calibration plot for the photosensor at the frame?

ps2frame.jpg

  5187   Thu Aug 11 11:50:56 2011 Manuel , IshwitaUpdatePEMCalibration of Guralp and STS2

 

 We just checked with a function generator the calibration of the ADC. We set a square wave with amplitude 1V. We measured the voltage with the oscilloscope and we found on the data viewer that one volt is 3208 counts. That's what we expected (+/- 10V for 16bits) but now we are more sure.

 

  5186   Thu Aug 11 10:56:08 2011 Ishwita, ManuelUpdatePEMMoving Seismometers

Quote:

We turned off the power of the seismometers and moved the Guralp1 close to the STS. Both are now situated below the center of the mode cleaner vacuum tube.

We oriented the X axis of the STS & Guralp1 along the X axis of the interferometer. Then we turned on the power again, but the STS channels don't give any signal. We think this is, because we didn't push the auto zero button.

 After pressing the auto-zero button (a lot of times) of the STS breakout box & aligning the bubble in the STS, we could finally get data from STS (Bacardi). So, now STS2 (Bacardi - Serial NR. 100151) is working!

  5185   Thu Aug 11 09:39:25 2011 Ishwita, ManuelUpdatePEMCalibration of Guralp and STS2

Quote:

Quote:

I'm pretty sure that don't have any ADC's with this gain. It should be +/- 10V for 16 bits. 

Jenne told us that the ADC was +/- 2V for 16 bits so our calibration is wrong. Since, the ADC is +/- 10V for 16 bits we need to change our calibration and now we can also use the purple STS breakout box.

 New calibration for Guralp:

ADC: 216counts = 20V Hence, calibration of ADC is (215x0.1) counts/V.


GURALP

Sensitivity = 800 V/ms-1

(215 x 0.1) counts/V x 800  V/ms-1 = 2621440 counts/ms-1 ----->  3.8147e-07 ms-1/count

Calibration = 3.8147e-07 ms-1/count


Using the above calibration we obtain the following plot:

pw_gurs_aug11.png

When we compare this plot with the old plot (see here) we see that in our calibration, we have got a factor of 10 less than the old plot. We do not know the gain of the Guralp. If we assume this missing 10 factor to be the gain of Guralp then we can get the same calibration as the old plot. But is it correct to do so?

  5184   Thu Aug 11 08:29:28 2011 steveUpdateComputersdataviewer at Rosalba

I'm having this problem with DTV every morning at Rosalba only. It wants to start with a negative GPS time and it can not connect to the frame builder.

Normally after a few time of starting it, it will work.

Attachment 1: gpsfmb.png
gpsfmb.png
  5183   Thu Aug 11 06:45:14 2011 NicoleSummarySUSShaking Testing

Koji and I have finished shaking the table for the first round of measurements (horizontal shaking). We have cleaned up the lab space used.

The FFT Analyzer has been put back to its position at the back side of the rack (near the seismometers).

 

I will calibrate the photosensor for the suspension frame and piece together/analyze/produce graphs of the data today. If everything is fine (the measurements are fine) and if there is a chance, we hope to shake the TT suspension vertically.

  5182   Thu Aug 11 04:45:07 2011 SureshUpdateIOOAligning the 1064nm beam with the in-vacuum pzt's

[Kiwamu, Suresh]

We worked on the beam path from MC to BS this evening. 

After the beam spots on MC1 and MC3 were close to the actuation nodes (<1mm away) we checked the beam position on the Faraday Isolator (FI) to make sure that it is passing through both the input and output apertures without clipping.  The beam is slightly displaced (by about half a beam diameter) downwards at the input of the FI.  The picture below is a screen shot from the MC1 monitor while Kiwamu held an IR card in front of the FI.

FI_input_spot.jpg

 

We then proceeded to check the beam position on various optical elements downstream.  But first we levelled the BS table and checked to see if the reflection from PJ1 (1st Piezo) is landing on the MMT1 properly.  It was and we did not make any adjustment to PJ1.  However the reflection from MMT1 was not centered on MMT2.  We adjusted the MMT1 to center the beam on it.  We then adjusted MMT2 to center the beam on PJ2.  At this point we noticed that the spot on IPPO (pick off window)  was off towards the right edge.  When we centered the beam on this it missed the center of the PRM.  In order to decide what needs to be moved, we adjusted PJ2 such that the beam hits the PR2, bounces back to PR3, and becomes co-incident with the green beam from X-arm on the BS.  Under this condition the beam is not in the center of PRM and nor is it centered on IPPO.  In fact it is being clipped at the edge of the IPPO. 

It is clear that both IPPO and the PRM need to be moved.  To be sure that the beam is centered on PR2 we plan to open the ITMX chamber tomorrow.

  5181   Thu Aug 11 02:16:57 2011 JenneUpdateGreen LockingY-green aligned and flashing

[Jenne, with ample supervision by Kiwamu and Suresh]

Y-green was aligned, and is now flashing.  The ETMY trans camera was very helpful for this alignment.  I didn't end up needing to use a foil aperture. 

Kiwamu and Suresh had just closed up the IOO doors, so we don't know yet where it's hitting on the PSL table (if the beam is making it that far).  Tomorrow we'll look at ITMY to see if the green beam is centered there, and if it's coming out to the PSL table.

  5180   Wed Aug 10 22:47:22 2011 ranaSummaryVACVacuum Workstation (linux3) re-activated

For some reason the workstation at the vac rack was off and unplugged. Nicole and I plugged its power back in to the EX rack.

I turned it on and it booted up fine; its not dead. To get it on to the network I just made the conversion from 131.215 to 192.168 that Joe had done on all the other computers several months ago.

Now it is showing the Vacuum overview screen correctly again and so Steve no longer has to monopolize one of the Martian laptops over there.

  5179   Wed Aug 10 20:40:17 2011 JennyUpdatePSLPDH locking: got an error signal

I ended up choosing a different dither frequency for driving the NPRO PZT: 230 kHz, because the phase modulation response in that region is higher according to other data taken on an NPRO laser (see this entry). At 230 there is a dip in the AM response of the PZT.

I am driving the PZT at 230 kHz and 13 dBm using a function generator. I am then monitoring the RF output of a PD that is detecting light reflected off the cavity. (The dither frequency was below the RF cutoff frequency of the PD, but it was appearing in the "DC output", so I am actually taking the "DC output" of the PD, which has my RF signal in it, blocking the real DC part of it with a DC block, and then mixing the signal with the 230kHz sine wave being sent to the PZT.

I am monitoring the mixer output on an oscilloscope, as well as the transmission through the cavity. I am sweeping the laser temperature using a lock in as a function generator sending out a sine wave at 0.2 V and 5 mHz. When there is a peak in the transmission, the error signal coming from the mixer passes through zero.

My next step is to find or build a low pass filter with a pole somewhere less than 100 kHz to cut out the unwanted higher frequency signal so that I have a demodulated error signal that I can use to lock the laser to the cavity.

 

  5178   Wed Aug 10 19:18:26 2011 NicoleSummarySUSFixed Reflective Photosensors; Recalibrated Photosensor 2

Thanks to Koji's help, the second photosensor, which was not working, has been fixed. I have re-calibrated the photosensor after fixing a problem with the circuit.  I have determined the new linear region to lie between 7.6 mm and 19.8mm. The slope defining the linear region is -0.26 V/mm (no longer the same as the first photosensor, which is -0.32 V/mm).

 

Here is the calibration plot.

ps2.jpg

  5177   Wed Aug 10 18:25:45 2011 JenneUpdateSUSETMY mini-update

[Jenne, Jamie]

ETMY is now in its new nominal position, according to the rails that Kiwamu put in the other day.  OSEM voltages are all centered, and the magnets looked pretty well centered in the OSEM bores.  We're taking data for some free swinging spectra, to check the decoupling. 

Next up: Align Y-green to the arm, then move on to fixing the other optics that Jamie pointed out.

  5176   Wed Aug 10 15:39:33 2011 jamieUpdateSUScurrent SUS input diagonalization overview

Below is the overview of all the core IFO suspension input diagonalizatidons.

Summary: ITMY, PRM, BS are really bad (in that order) and are our top priorities.

UPDATE:

I had originally put the condition number of the calculated input matrix (M) in the last column.  However, after some discussion we decided that this is not in fact what we want to look at.  The condition number of a matrix is unity if the matrix is completely diagonal.  However, even our ideal input matrix is not diagonal, so the "best" condition number for the input matrix is unclear.

What instead we do know is that the matrix, B, that describes the difference between the calculated input matrix, M, and the ideal input matrix, M0: should be diagonal (identity, in fact):

M = M0 B

B should be diagonal (identity, in fact), and it's condition number should ideally be 1.  So now we calculate B-1, since it can be calculated from the pre-inverted input matrix:

B-1 = M-1 * M0

From that we calculate cond(B) == cond(B-1).

cond(B) is our new measure of the "badness" of the OSEMS.

new summary: ITMY, PRM, BS are really bad (in that order) and are our top priorities.

 

TM    INPUT MATRIX (M)
 cond(M) cond(B)
 PRM PRM.png        pit     yaw     pos     side    butt
UL   -2.000  -2.000  -2.000  -0.345   2.097 
UR   -0.375  -0.227  -0.312  -0.060   0.247 
LR    1.060   1.075   0.971   0.143  -0.984 
LL   -0.565  -0.698  -0.717  -0.141   0.672 
SD    1.513   1.485   1.498   1.000  -1.590
 75.569 106.756
 SRM SRM.png  

      pit     yaw     pos     side    butt
UL    0.791   1.060   1.114  -0.133   1.026 
UR    1.022  -0.940   1.052  -0.061  -1.027 
LR   -0.978  -0.987   0.886  -0.031   0.903 
LL   -1.209   1.013   0.948  -0.103  -1.043 
SD    0.286   0.105   1.249   1.000   0.030

 2.6501 3.90776
 BS  BS.png  

      pit     yaw     pos     side    butt
UL    1.420   0.818  -0.069   0.352   1.038 
UR    0.276  -1.182   1.931  -0.217  -0.905 
LR   -1.724  -0.274   1.940  -0.254   0.862 
LL   -0.580   1.726  -0.060   0.315  -1.194 
SD    0.560   0.171  -3.535   1.000   0.075 

9.8152 7.28516
ITMX ITMX.png       pit     yaw     pos     side    butt
UL    0.437   1.015   1.050  -0.065   0.714 
UR    0.827  -0.985   1.129  -0.221  -0.957 
LR   -1.173  -1.205   0.950  -0.281   1.245 
LL   -1.563   0.795   0.871  -0.125  -1.084 
SD   -0.581  -0.851   2.573   1.000  -0.171 
 4.08172 4.69811
 ITMY  ITMY.png  

      pit     yaw     pos     side    butt
UL    0.905  -0.884  -0.873   0.197   0.891 
UR   -1.095   1.088   1.127  -0.252  -1.115 
LR   -0.012  -0.028   0.002   0.001   0.030 
LL    1.988  -2.000  -1.998   0.451   1.964 
SD    4.542  -4.608  -4.621   1.000   4.517 

 801.453 774.901
 ETMX  ETMX.png        pit     yaw     pos     side    butt
UL    0.344   0.475   1.601   0.314   1.043 
UR    0.283  -1.525   1.786  -0.071  -1.181 
LR   -1.717  -1.569   0.399  -0.102   0.938 
LL   -1.656   0.431   0.214   0.283  -0.837 
SD    0.995  -2.632  -0.999   1.000  -0.110 
 4.26181 4.33518
 ETMY  ETMY.png        pit     yaw     pos     side    butt
UL   -0.212   1.272   1.401  -0.127   0.941 
UR    0.835  -0.728   1.534  -0.101  -1.054 
LR   -0.953  -1.183   0.599  -0.066   0.827 
LL   -2.000   0.817   0.466  -0.092  -1.177 
SD   -0.172   0.438   2.238   1.000  -0.008 
 4.04847 4.33725

 

  5175   Wed Aug 10 15:17:39 2011 Ishwita, ManuelUpdatePEMCalibration of Guralp and STS2

Quote:

I'm pretty sure that don't have any ADC's with this gain. It should be +/- 10V for 16 bits. 

Jenne told us that the ADC was +/- 2V for 16 bits so our calibration is wrong. Since, the ADC is +/- 10V for 16 bits we need to change our calibration and now we can also use the purple STS breakout box.

  5174   Wed Aug 10 14:35:39 2011 ranaUpdatePEMCalibration of Guralp and STS2

I'm pretty sure that don't have any ADC's with this gain. It should be +/- 10V for 16 bits. 

  5173   Wed Aug 10 14:30:55 2011 kiwamuUpdateIOORe: MC A2L alignment

I modified a set of the automated MC locking scripts which are dedicated for the low power MC.

Currently there are three scripts like the usual MC locking scripts:

(1)mcup_low_power, (2) mcdown_low_power and (3) autolockMCmain40_low_power.

I ran those scripts on op340m as usual and so far they are running very well. The lock acquisition is quite repeatable.

I hope theses scripts always bring the lock condition to the same one and hence the LOCKIN signals don't change by every lock.

 

- To run the script

  log into op340m and run autolockMCmain40m_low_power

Quote from #5167

And the MC settles into a new position when the MC-PSL servo loop was disturbed by random denizens in the lab.  Requiring us to start over again.

 

  5172   Wed Aug 10 14:27:39 2011 Ishwita , ManuelUpdatePEMCalibration of Guralp and STS2

ADC: 216counts = 4V Hence, calibration of ADC is 214counts/V.

Gain of the AA board, g1 = 0.1


GURALP

Sensitivity = 800 V/ms-1

214 counts/V x 800  V/ms-1 = 13107200 counts/ms-1 -----> 7.6294e-08 ms-1/count

Gain, g2 = 10

Calibration = 7.6294e-08 ms-1/count x g1 x g2 = 7.6294e-08 ms-1/count


 STS

Sensitivity = 1500 V/ms-1

214 counts/V x 1500  V/ms-1 = 24576000 counts/ms-1 -----> 4.069e-08 ms-1/count

Gain of the STS electronic breakout box, g3 = 10

Calibration = 4.069e-08 ms-1/count x g1 x g3 = 4.069e-08 ms-1/count

  5171   Wed Aug 10 13:52:23 2011 Manuel , IshwitaUpdatePEMMoving Seismometers

We turned off the power of the seismometers and moved the Guralp1 close to the STS. Both are now situated below the center of the mode cleaner vacuum tube.

We oriented the X axis of the STS & Guralp1 along the X axis of the interferometer. Then we turned on the power again, but the STS channels don't give any signal. We think this is, because we didn't push the auto zero button.

  5170   Wed Aug 10 12:33:34 2011 Manuel, IshwitaSummaryPEMWeekly summary

We got the results of the wiener filtering simulations (Elog Entry)

We got the power spectra and coherence of the seismic noise measurements from GURALPs and STS seismometers (Elog Entry)

We tried to whiten the target and the input signal for the computation of the wiener filter for the real data, but the results are unsatisfactory. We should not care about high frequencies in wiener filter computation so we will just filter them off in the filter output with a low pass filter.

We just found the right gain for the system seismometer-AAboard-ADC (Elog Entry)

  5169   Wed Aug 10 12:32:09 2011 NicoleSummarySUSWeekly Summary Update

Last night, I attached a metal plate to the Vout faceplate of my photosensor circuit box because the BNC connection terminals were loose. This was Jamie's suggestion to establish a more secure connection (I had originally drilled holes for the BNCs that were much too large).

 

I have also fixed the mechancial set-up of my shaking experiment so that the horizontal sliding platform does not interfere with the photodiode mounting stage. Koji pointed out last night that in the full range of motion, the photodiode mounting stage interferes with the movement of the sliding platform when the platform is at its full range.

 

I have began shaking. I am getting a problem, as my voltage outputs are just appearing a high-frequency noise.

  5168   Wed Aug 10 12:28:22 2011 Ishwita , ManuelUpdatePEMAA board gain

We used a function generator, an oscilloscope and the Data Viewer to check the gain of the new AA board (used for the seismometers). Putting a sine wave of 0.3V (using a function generator) to the AA board, we could see about 500 counts in the Data Viewer. The calibration of the ADC is 214 counts/volt, so the AA board gives to the ADC an output of 0.03V. This proves that the AA board has a gain of 0.1. Guralp1 and STS1 (Bacardi), both have a gain of 10 now, that balance the AAboard gain of 0.1. If we consider the gain of AA board in our calibrated power spectrum plot of seismic signals from Guralp1 and STS1 (Bacardi), we get the following plot:

pw_gur1_sts1_aug10.png

  5167   Wed Aug 10 11:22:55 2011 SureshUpdateIOOMC A2L alignment

[Kiwamu, Suresh]

We attempted to minimise the A2L coupling in the MC mirrors (centering the beam spot on the actuation node on each optic).  While it was easy to minimise the coupling in the pitch for all the three optics and yaw of MC2, the yaw alignment of MC1 and MC3 proved to be difficult.  For one the adjustment required was quite large, so much so that PSL alignment into the MC is often lost during this adujstment.  We had to align the PSL coupling several times in order to proceed.   And the MC settles into a new position when the MC-PSL servo loop was disturbed by random denizens in the lab.  Requiring us to start over again.

Kiwamu wrote a script to measure the MC(optic)(Pitch/yaw) -> Lockin(#1 to #6) matrix.  Inverting this matrix gave us the linear combination of the offsets to put on the MC# PIT/YAW  inorder to minimise a specific Lockin output.  However the cross couplings were not completely eliminated.  This made it very hard to predict what a given set of offsets were going to do to the Lockin outputs.

Net result:  the spots are centered in vertical direction (pitch) but not in the horizontal (yaw)

Day time tasks have started so I am quitting now.

  5166   Wed Aug 10 07:59:33 2011 steveUpdateSUSthis is too dam nice

Quote:

Suresh and I tweaked the OSEM angles in ETMX yesterday.  Last night the ETMs were left free swinging, and today I ran Rana's peakFit scripts on ETMX to check the input diagnolization:

ETMX.png

It's well inverted, but the matrix elements are not great:

       pit       yaw       pos       side      butt
UL    0.3466    0.4685    1.6092    0.3107    1.0428
UR    0.2630   -1.5315    1.7894   -0.0706   -1.1859
LR   -1.7370   -1.5681    0.3908   -0.0964    0.9392
LL   -1.6534    0.4319    0.2106    0.2849   -0.8320
SD    1.0834   -2.6676   -0.9920    1.0000   -0.1101

The magnets are all pretty well centered in the OSEMS, and we worked at rotating the OSEMS such that the bounce mode was minimized.

Rana and Koji are working on ETMY now.  Maybe they'll come up with a better procedure.

 

  5165   Wed Aug 10 02:40:40 2011 JennyUpdatePSLDither freq for PZT chosen: 2.418 MHz

I've finished using the network analyzer to characterize find a dither frequency for driving the PZT to use in my PDH locking. I found a region in which the amplitude response of the PZT is low: The dip is centered at 2.418 MHz. Changing the NPRO laser temperature by 100mK has no significant effect on the transfer function in that region. I will post plots tomorrow.

I'm finished with the network analyzer. It is unplugged, and the cart is still near the PSL table. (I'll roll it back tomorrow when it won't disturb interferometer locking).

I closed the shutter on the NPRO at the end of the night.

Tomorrow I plan to put together the fast locking setup. I'll drive the PZT at 2.418 MHz. More details to come tomorrow.

  5164   Wed Aug 10 02:29:38 2011 JenneUpdateSUSETMY exploration

Quote:

- Jenne will make a better kick/free-swing test later.

 02:27am, ran the new freeswinging-ifo.csh script.  It's just a copy of freeswinging-all.csh, but it doesn't include the MC mirrors, since Suresh and Kiwamu are still working.  

Now we have copies of the script for -all, -mc, -ifo to cover the various sections of the suspended interferometer.

  5163   Wed Aug 10 01:40:40 2011 KojiUpdateSUSETMY exploration

[Rana Koji Jenne Jamie]

- The situation of the ETMY suspension is improved.
- The damping servos except for Pitch are now functional.
- We intentionally turned off the damping servos for the matrix measurements.


- Opened the light door of the ETMY chamber.

- We set up the CDS SUS lockin:

        Excite UL/UR/LL/LR equally [by setting the output matrix (1, 1, 1, 1, 0)] at 3.12Hz with 2000 cnts
        Put the OSEM PD outputs into lockin one by one. For the image rejection, 0.1Hz 4th order LPF has been used though we like to use a faster settling LPF.

- Found only UL was responding to the excitation. After fitzing with the cables and connectors, it was found that the DAC card was loose from the bus.
  By pushing the card the responses have been back. [Note we had the reboot of c1iscey almost at the same time.]

- Checked the response in the I channel of the lockin.
        UL -8ish / UR +7ish / LR +5ish / LL +2ish

- Tweaked LL sensor to get better response ==> in vain. Decided to move the lower OSEM plate for the better positioning of the LR/LL.
- Got reasonable (+5ish) response for LL.

- Confirmed that the POS/YAW/SIDE damping works with positive gains. PITCH did not work with the negative gain (but that could be a good sign.)

- Let the suspension freely swinging for a while (~30min). Checked the side/pos separation. They are not perfect but seemed diagonalizable.

- Closed the light door.

- Jenne will make a better kick/free-swing test later.

  5162   Wed Aug 10 00:21:10 2011 jamieUpdateCDSupdates to peakFit scripts

I updated the peakFit routines to make them a bit more user friendly:

  • modified so that any subset of optics can be processed at a time, instead of just all
  • broke out tweakable fit parameters into a separate parameters.m file
  • added a README that describes use

These changes were committed to the 40m svn.

  5161   Wed Aug 10 00:11:39 2011 jamieUpdateSUScheck of input diagnolization of ETMX after OSEM tweaking

Suresh and I tweaked the OSEM angles in ETMX yesterday.  Last night the ETMs were left free swinging, and today I ran Rana's peakFit scripts on ETMX to check the input diagnolization:

ETMX.png

It's well inverted, but the matrix elements are not great:

       pit       yaw       pos       side      butt
UL    0.3466    0.4685    1.6092    0.3107    1.0428
UR    0.2630   -1.5315    1.7894   -0.0706   -1.1859
LR   -1.7370   -1.5681    0.3908   -0.0964    0.9392
LL   -1.6534    0.4319    0.2106    0.2849   -0.8320
SD    1.0834   -2.6676   -0.9920    1.0000   -0.1101

The magnets are all pretty well centered in the OSEMS, and we worked at rotating the OSEMS such that the bounce mode was minimized.

Rana and Koji are working on ETMY now.  Maybe they'll come up with a better procedure.

  5160   Tue Aug 9 19:53:56 2011 NicoleSummarySUSWeekly Summary

This week, I have finished assembling everything I need to begin shaking. I built an intermediary mounting stage to mount the TT suspension base to the horizontal sliding platform, finished assembling the second photodiode, finished assembling the photosensor circuit box, and calibrated the two photosensors. Today I built a platform/stage to mount the photodiodes so that they are located close enough to the mirror/suspension that they can operate in the linear range.  Below is an image of the set-up.

entiresetup.jpg

The amplifer that Koji fixed is acting a bit strange again...It is sometimes shutting off (Apparently, it can only manage to do short runs ~ 1minute? That should be enough time?).

The set-up is ready to begin taking measurements.

  5159   Tue Aug 9 16:54:47 2011 steveUpdateVACscratch on vac door

Jenne found a big scratch on the north vac door of ITMY. Fortunately it does not reach the inner annulos 0 -ring seal.

This is precisely what we have to avoid to preserve our vacuum system!

 

Attachment 1: P1080152.JPG
P1080152.JPG
ELOG V3.1.3-